# SPRec: Leveraging Self-Play to Debias Preference Alignment for Large Language Model-based Recommendations

### v1

@@ 5.3.2. SPRec without DPO
Removing DPO reduces SPRec to further SFT training on new data, ensuring that performance gains are not due to incorporating additional data. As shown in Table 3, even with new data, additional SFT fails to improve diversity or fairness metrics, and the recommendations remain biased. Furthermore, Fig. 4 reveals minimal fluctuations during training, indicating that the prior SFT training has already converged. This experiment underscores the limitations of SFT-based methods in addressing recommendation fairness and diversity.
@@ 5.3.3. Randomly sampling negative items
Replacing the negative sampling strategy with random sampling (SPRec-RN) fails to improve DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets (Table 3). Additionally, SPRec-RN exhibits worse fairness metrics compared to SPRec. While it achieves a significant boost in DivRatio on the CDs dataset, its ORRatio remains poor. These results suggest that random negative sampling is ineffective in suppressing popular items, leaving a pronounced long-tail effect in the recommendations. This experiment demonstrates that our self-play negative sampling strategy effectively balances the model’s output distribution, resulting in debiased recommendations.
@@ 5.4. Impact of Negative Samples (RQ3)
We investigate the role of negative samples in SPRec by introducing a proportion of random negative samples to contaminate SPRec’s original self-play samples. Additionally, we examine the impact of increasing the number of negative samples in SPRec’s loss function (Eq. (8)). To achieve this, we adopt the SDPO loss function to expand to multiple negative samples. For efficiency, we limit the training sample size to 1,024, keeping other experimental settings consistent with Section 5.2. To ensure result stability, the model’s performance is averaged over the last three training iterations.
@@ 5.4.2. Effect of Negative Sample Size
6. Conclusion
This paper theoretically and experimentally analyzes the limitations of existing SFT- and DPO-based methods in LLM-based recommender systems. These approaches often promote a narrow set of items, intensifying the filter bubble effect, harming user interests, and raising fairness concerns that may undermine creators’ motivation. To address these challenges, we propose SPRec, a novel self-play recommendation framework that mitigates over-recommendation and enhances fairness without requiring additional data or manual intervention. SPRec leverages the model’s output from the previous round as negative samples, effectively re-weighting the DPO loss function with the model’s logits. This adaptive mechanism suppresses biased items and improves recommendation quality.
We conducted experiments on four real-world datasets, validating the effectiveness of our method in improving recommendation accuracy, diversity, and fairness. Comprehensive ablation studies revealed that fairness improvements largely stem from the suppression of self-play negative samples during DPO. In the future, we plan to explore more advanced preference alignment techniques for LRSs, aiming for smarter personalized decision-making and more helpful and harmless recommendation systems.
@@ Appendix A Mathematical Derivations
The DPO loss is derived from the RLHF objective as follows:
@@
In the original paper of DPO (Rafailov et al., 2024), the authors proved that the optimal policy $\pi_{\theta}^{*}$ for DPO loss in Eq. (LABEL:eqn:dpo_def) and the solution to the optimization problem in Eq. (9) are the same. Thus, we can analyze the solution to Eq. (9), equivalent to examining the DPO loss.
@@
which has a well-known closed form solution (Rafailov et al., 2024): π_θ^*(y—x) ∝π_ref(y—x)⋅(pD(y—x)qD(y—x))^1/β. ∎
Appendix B Implementation Details
For all traditional models, we set the sizes of the training and validation sets to $65,536$ and $8,192$, respectively. The embedding size was fixed at $64$, and the dropout ratio was set to 0.1. Negative samples were randomly sampled in training, with Adam as the optimizer and a learning rate of 4e-3. For LLM-based methods, we adopted Llama-3.2-1B-Instruct as the backbone LLM. Considering the ability of LLMs to quickly adapt to downstream tasks with limited data, we followed BIGRec (Bao et al., 2023a) and used relatively smaller datasets compared to traditional models. For SFT-based methods, including BIGRec and Re-weighting, the sizes of the training and validation sets were set to $4,096$ and $512$, respectively. For DPO-based methods, the sizes were set to $2,048$ and $256$, and training was uniformly conducted on the basis of BIGRec. For SPRec, the total number of iterations was set to $5$, with each SFT and DPO phase trained for 1 epoch. All experiments were carried out on four RTX 3090 GPUs, each with 24GB of VRAM.

### v3

@@ 6.3.2. SPRec without DPO
Removing DPO reduces SPRec to further SFT training, ensuring that performance gains are not due to incorporating additional data. As shown in Table 2, additional SFT fails to improve diversity or fairness metrics, and the recommendations remain biased. Furthermore, Fig. 5 reveals minimal fluctuations during training, indicating that the prior SFT training has already converged. This experiment underscores the limitations of SFT-based methods in addressing recommendation fairness and diversity.
@@ 6.3.3. Randomly sampling negative items
As observed in Table 2, when the negative sampling strategy is replaced with random sampling, SPRec-RN fails to achieve further improvements in DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets. Additionally, SPRec-RN’s fairness metrics perform worse compared to SPRec. Although SPRec-RN shows a significant improvement in DivRatio on the CDs and Vinyl dataset, its ORRatio still performs poorly. This suggests that random sampling of negative samples during training is ineffective at suppressing popular items, and the recommendation results continue to exhibit a significant long-tail effect. This ablation experiment demonstrates that our Self-play negative sampling strategy effectively balances the distribution of the model’s output, leading to debiasing in recommendations. Replacing the negative sampling strategy with random sampling (SPRec-RN) fails to improve DivRatio and ORRatio metrics on the MovieLens and Goodreads datasets (Table 2). Additionally, SPRec-RN exhibits worse fairness metrics compared to SPRec. While it achieves a significant boost in DivRatio on the CDs dataset, its ORRatio remains poor. These results suggest that random negative sampling is ineffective in suppressing popular items, leaving a pronounced long-tail effect in the recommendations. This experiment demonstrates that our self-play negative sampling strategy effectively balances the model’s output distribution, resulting in debiased recommendations.
@@ 6.4. Impact of Negative Samples (RQ3)
We investigate the role of negative samples in SPRec by introducing a proportion of random negative samples to contaminate SPRec’s original self-play samples. Additionally, we examine the impact of increasing the number of negative samples in SPRec’s loss function (Eq. (6)). To achieve this, we adopt the SDPO loss function to expand to multiple negative samples. For efficiency, we limit the training sample size to 1,024, keeping other experimental settings consistent with Section 6.2. To ensure result stability, the model’s performance is averaged over the last three training iterations. We report the results on the Movielens dataset.
@@ 6.4.2. Effect of Negative Sample Size
7. Conclusion & Discussion
Our work establishes a critical bridge between preference alignment techniques and fairness-aware recommendation in the era of LLMs. Through both theoretical analysis and empirical validation, we demonstrate that conventional DPO-based tuning fundamentally conflicts with the principles of equitable recommendation, creating self-reinforcing popularity biases that traditional debiasing approaches fail to address. The proposed SPRec framework represents a paradigm shift in recommendation alignment - rather than treating bias mitigation as a post-hoc correction, we redesign the core learning mechanism to enable autonomous bias suppression through self-regulated competition between model generations. This approach not only achieves state-of-the-art performance across accuracy and fairness metrics but more importantly, provides a blueprint for developing self-correcting AI systems that maintain alignment with both user preferences and ethical constraints.
Despite its effectiveness, our work primarily addresses bias in DPO-based tuning, while overlooking the popularity bias already present in SFT due to its cross-entropy loss. Future research should focus on mitigating bias at the SFT stage to ensure fairness from the start of fine-tuning. Additionally, optimizing preferences in recommendation is a long-term challenge, requiring alignment across sequential recommendations rather than individual predictions. However, LLMs generate outputs token by token, making it difficult to optimize preferences from token-level to item-level and ultimately list-level recommendations. Tackling this issue will require new datasets, benchmarks, and models capable of long-term alignment. A promising direction is reinforcement learning with process-level rewards, shifting optimization from short-term token likelihood to long-horizon user engagement.
Acknowledgements
This work is supported by the National Key Research and Development Program of China (2021ZD0111802), the National Natural Science Foundation of China (62402470, 62272437, U24B20180, 62121002), the Fundamental Research Funds for the Central Universities of China (WK2100000053, PA2024GDSK0107), Anhui Provincial Natural Science Foundation (2408085QF189), and the Postdoctoral Fellowship Program of CPSF (GZC20241643). This research is supported by the advanced computing resources provided by the Supercomputing Center of the USTC.
@@ Appendix A Mathematical Derivations
The DPO loss is derived from the objective of Reinforcement Learning with Human Feedback (RLHF):
@@
In the original paper of DPO (Rafailov et al., 2024), the authors proved that the optimal policy $\pi_{\theta}^{*}$ for DPO loss in Eq. Eq. (4) and the solution to the optimization problem in Eq. (7) are the same. Thus, we can analyze the solution to Eq. (7), equivalent to examining the DPO loss.
@@
which has a well-known closed form solution (Rafailov et al., 2024): π_θ^*(y—x) ∝π_ref(y—x)⋅(pD(y—x)qD(y—x))^1/β.
Appendix B Dataset Statistics
Our datasets span diverse domains, including movies, books, music, and games, offering varied sizes and user interaction patterns to provide a comprehensive basis for evaluating LRSs. Note that while we report the full dataset statistics, only a subset of interaction sequences is sampled for LLM fine-tuning, as detailed in Section 6.1.1.
Table 3. Statistics of Datasets.
Datasets #Items #Interactions #Sequences
MovieLens 10,682 10,000,054 9,301,274
Goodreads 4,058 160,398 6,031
CDs_and_Vinyl 13,078 185,855 21,347
Steam 32,094 178,961 29,876
Appendix C Implementation Details
For LLM-based methods, we adopted Llama-3.2-1B-Instruct as the backbone LLM. Considering the ability of LLMs to quickly adapt to downstream tasks with limited data, we followed BIGRec (Bao et al., 2025) and used relatively smaller datasets. To ensure fairness in comparison, all baseline methods and SPRec utilize the same dataset as used in the SFT training phase. For SPRec, the total number of iterations was set to $5$, with each SFT and DPO phase trained for 1 epoch. To ensure that the training data used in each iteration is not identical, we further randomly sample half of the training data (i.e., 2048 interactions) for training in each iteration. All experiments were carried out on four RTX 3090 GPUs, each with 24GB of VRAM.
For the traditional model SASRec, we use the same training and validation datasets as other LLM-based methods, with dataset sizes of 4,096 and 512, respectively. The embedding size was fixed at $64$, and the dropout ratio was set to 0.1. Negative samples were randomly sampled in training, with Adam as the optimizer and a learning rate of 4e-3. More details of the implementation are available via https://github.com/RegionCh/SPRec.